OpenStudy (anonymous):

The price of a notebook has risen to $3.35 today. Yesturday's price was $3.10. find the percentage increase

hero (hero):

\(\%\Delta = \frac{|a-b|}{a} \times 100\) a = yesterday's price b = today's price

OpenStudy (anonymous):

thanks so much dude